Pablo needs to produce 10 liters of a 15% soultion from 10% to 25% solutions. How many liters of 10% solution does he need?
Answer by jorel1380(3719)

10x.15=1.5
m+n=10
.1m+.25n=1.5
10m+25n=150
10m+10n=100
15n=50
n=50/15=3.333
m=6.66666
Pablo needs 6.666 liters of 10% solution